# Rethinking Negative Instances for Generative Named Entity Recognition

### v1

@@ 5.1 Settings
Our main point of comparison is UniversalNER Zhou et al. (2023) as it is the approach closest to our system, with similar data and training procedures. Another baseline considered for comparison is GLiNER Zaratiana et al. (2023), which utilizes bi-directional models to match entity types with textual spans in a latent space. We also include some strong Information Extraction (IE) systems like InstructUIE Wang et al. (2023), which is based on Flan-T5-xxl Chung et al. (2022) and fine-tuned on diverse information extraction datasets, and GoLLIE Sainz et al. (2023), which is based on Code-LLaMA Roziere et al. (2023), and use guidelines to improve model’s zero-shot performance. We use strict entity-level micro-$F_{1}$ as the evaluation metric for comparison.
@@
Table 4: Model’s performance and inference speed in zero-shot and supervised settings. The inference speed is tested in a single A100 node with batch size 4 per device. More details are outlined in Appendix E.
@@ 5.3 Supervised Evaluation
Table 5: Ablation study of Hierarchical Matching.
5.4 Ablation Results
We have demonstrated the effectiveness of negative instances in our preliminary study. In this part, we conduct the ablation study in Table 5 to evaluate the performance of our Hierarchical Matching Algorithm. Removing Back Tokenization and the LCS algorithm results in a decrease in effectiveness, underscoring the efficacy of our Hierarchical Matching algorithm.
@@ 6 Analysis
Beam search can enhance the performance of generative models by expanding the search space to include multiple hypotheses at each generation step. Previous research Yan et al. (2021) has demonstrated that applying beam search in an entity-centric generation does not improve the model’s performance or even degrade it. We conduct experiments on UniNER and our model, the results of which are shown in Table 6. We discover that as the beam size increases, the performance of the UniNER model decreases. In contrast, we observe a consistent improvement with beam search under our task schema. Upon a detailed case study of the model’s generated results, we found that our task schema possesses a self-correction mechanism. The model retains some other hypotheses while generating subsequent results. In the decoding process that follows, the model can correct earlier mistakes. As demonstrated in Fig. 6, the model revises its previous incorrect prediction of “O” for “the fog” upon encountering the subsequent token “rated”. This token is crucial for identifying the entity type associated with “the fog”. A detailed analysis is provided in Appendix D.
@@ 7 Conclusion
This paper explores the potential of a strong Generative Named Entity Recognition (NER) system based on pre-trained LLMs by integrating the negative instances into training. Through experiments, we have demonstrated significant advancements. Our approach, which combines the inclusion of contextual information and a clear definition of entity boundaries through negative instances, has proven to be highly effective in improving the model’s performance, especially in zero-shot scenarios where prediction uncertainty is high. The introduction of a Hierarchical Matching algorithm further addresses the challenges of converting unstructured text into structured entities, ensuring accurate categorization and alignment. These findings highlight the crucial role of negative instances in NER tasks and the potential of generative models to revolutionize the field.
@@ Appendix B Hyper-parameters settings
GNER LLaMA Omission (39%) 1 raw
who directed the film the lorax
pred.
who directed the lorax
2 raw
any reasonably priced indian restaurants in the theater district
pred.
any reasonably priced indian restaurants in theater district
Addition (3%) 3 raw
the conservative regionalist navarra suma finished first and  …
pred.
the conservative regionalist regionalist navarra suma finished first and  …
Substitution (58%) 4 raw
which five star italian restaurants in manattan have the best reviews
pred.
which five star italian restaurants in manhattan have the best reviews
5 raw
polyethylene terephthalate ( pet ) bottles are made from ethylene and p-xylene .
pred.
polyethylene terephthalate ( p e t) bottles are made from ethylene and p-xylene .
GNER T5 Omission (23%) 6 raw
…  whose debut album tol cormpt norz norz norz rock hard journalist wolf-rüdiger mühlmann considers a part of war metal ’s roots .
pred.
…  whose debut album tol cormpt norz norz rock hard journalist wolf-rüdiger mühlmann considers a part of war metal ’s roots .
7 raw
jennifer lien starred in this action film of the the last six years that received a really good rating
pred.
jennifer lien starred in this action film of the last six years that received a really good rating
Addition (2%) 8 raw
…  performed by wet wet wet that remained at number 1  …
pred.
…  performed by wet wet wet wet that remained at number 1  …
9 raw
…  liked by many people that starred william forsythe
pred.
…  liked by many people that starred william forsythe the
Substitution (75%) 10 raw
four more children followed : charlotte brontë , ( 1816-1855 ) , branwell brontë ( 1817-1848 ) , emily brontë , ( 1818-1848 ) and anne ( 1820-1849 ) .
pred.
four more children followed : charlotte bront , ( 1816-1855 ) , branwell bront ( 1817-1848 ) , emily bront , ( 1818-1848 ) and anne ( 1820-1849 ) .
@@ Appendix C Problems in long sequence
Appendix D Self-correction Mechanism
@@
Appendix E Detailed Evaluation Results
@@
GNER LLaMA 1 w/o beam search
who(O) is(O) directing(O) the(O) hobbit(B-title)
w/ beam search
who(O) is(O) directing(O) the(B-title) hobbit(I-title)
2 w/o beam search
what(O) is(O) the(O) plot(O) of(O) the(O) wild(B-title) bunch(I-title)
w/ beam search
what(O) is(O) the(O) plot(O) of(O) the(B-title) wild(I-title) bunch(I-title)
3 w/o beam search
was(O) there(O) a(O) romantic(O) film(O) noir(O)
w/ beam search
was(O) there(O) a(O) romantic(B-genre) film(I-genre) noir(I-genre)
4 w/o beam search
does(O) paymon(B-Restaurant Name) serves(O) white(B-Cuisine) wine(I-Cuisine)
w/ beam search
does(O) paymon(B-Restaurant Name) serves(O) white(B-Dish) wine(I-Dish)
GNER T5 5 w/o beam search
…  some(O) batman(B-character) movies(O) from(O) the(O) 1990s(B-year)
w/ beam search
…  some(O) batman(B-title) movies(I-title) from(O) the(O) 1990s(B-year)
6 w/o beam search
where(O) was(O) the(O) presidio(B-title) filmed(O)
w/ beam search
where(O) was(O) the(B-title) presidio(I-title) filmed(O)
7 w/o beam search
…  the(O) third(O) harry(O) potter(O) movie(O) called(O)
w/ beam search
…  the(O) third(B-title) harry(I-title) potter(I-title) movie(I-title) called(O)
@@
Appendix F Environmental Impact

### v2

@@ 5 Experiments
ChatGPT - 52.4 39.8 66.6 68.5 67.0 5.3 32.8 47.5
InstructUIE flan-t5-xxl (11B) 48.4 48.8 54.4 49.9 49.4 63.0 21.0 47.8
GoLLIE-7B Code-LLaMA-7B 60.3 67.1 64.5 60.8 60.5 63.0 43.4 59.9
GoLLIE-13B Code-LLaMA-13B 63.8 60.1 68.5 56.2 61.5 62.5 49.8 60.3
UniNER-7B LLaMA-7B 53.5 59.4 65.0 60.8 61.1 42.4 31.7 53.4
UniNER-13B LLaMA-13B 54.2 60.9 64.5 61.4 63.5 48.7 36.2 55.6
GLiNER-L DeBERTa-v3-300M 60.6 68.4 69.5 74.8 69.4 57.2 42.8 63.2
GNER-T5 flan-t5-base (250M) 56.8 58.7 72.3 64.5 68.0 54.5 41.4 59.5
flan-t5-large (780M) 62.6 58.2 76.7 67.0 72.6 58.6 48.6 63.5
flan-t5-xl (3B) 62.1 64.9 80.6 73.7 68.7 63.0 49.8 66.1
flan-t5-xxl (11B) 68.2 68.7 81.2 75.1 76.7 62.5 51.0 69.1
GNER-LLaMA LLaMA-7B 63.1 68.2 75.7 69.4 69.9 68.6 47.5 66.1
Table 2: Zero-shot evaluation results, where $\dagger$ denotes IE Models and $\ddagger$ denotes NER Models. Results for ChatGPT and UniNER are from Zhou et al. (2023); InstructUIE are from Wang et al. (2023); GoLLIE are from Sainz et al. (2023); GLiNER-L are from Zaratiana et al. (2023). We bold the best results and underline the second-best results. More details about the performance including error bars are shown in Appendix F.
@@ 5.1 Settings
Our main point of comparison is UniversalNER Zhou et al. (2023) as it is the approach closest to our system, with similar data and training procedures. Another baseline considered for comparison is GLiNER Zaratiana et al. (2023), which utilizes bi-directional models to match entity types with textual spans in a latent space. We also include some strong Information Extraction (IE) systems like InstructUIE Wang et al. (2023), which is based on Flan-T5-xxl Chung et al. (2022) and fine-tuned on diverse information extraction datasets, and GoLLIE Sainz et al. (2023), which is based on Code-LLaMA Roziere et al. (2023), and use guidelines to improve model’s zero-shot performance. We use strict entity-level micro-$F_{1}$ as the evaluation metric for comparison. Previous work lack a uniform setting; UniNER removed the “else” entity type in the CrossNER dataset, while InstructUIE, GoLLIE, and GLiNER retained it. To standardize the settings, we reevaluated their methods using their publicly released checkpoints and code, ensuring that all test set configurations are consistent with UniNER.
@@
Table 4: Model’s performance and inference speed in zero-shot and supervised settings. The inference speed is tested in a single A100 node with batch size 4 per device. More details are outlined in Appendix F.
@@ 5.3 Supervised Evaluation
Sequence Length 0-60 60-100 100-200
LCS $O(N^{2})$ 1.0$\times$ 1.0$\times$ 1.0$\times$
LCS $O(N\log N)$ 2.9$\times$ 5.5$\times$ 6.1$\times$
LCS (ours) 3.8$\times$ 12.6$\times$ 17.3$\times$
Table 5: Acceleration effect of our optimized algorithm across different length ranges. All samples are selected from the generated results under supervised settings.
5.4 Ablation Results
We have demonstrated the effectiveness of negative instances in section 3. In this part, we conduct the ablation study to evaluate the performance of our LCS Matching Algorithm. Our focus lies in two aspects: (1) how optimization in the algorithm increases the model’s inference efficiency, and (2) how optimization in the matching condition improves the model’s performance. For the former, we evaluate the algorithm’s acceleration across various sentence length ranges, as shown in Table 5. As sentence length increases, the acceleration effect of our algorithm becomes more pronounced. For sentence lengths between 100 and 200, it achieves an average acceleration factor of 17.3. For the latter, we remove the Back Tokenization procedure from LCS and eliminate all LCS processes. The results, presented in Table 6, indicate that removing Back Tokenization and the whole LCS algorithm leads to a decrease in effectiveness, underscoring the efficacy of our LCS Matching algorithm.
@@ 5.4 Ablation Results
Table 6: Ablation study of LCS Matching and Back Tokenization.
@@ 6 Analysis
Beam search can enhance the performance of generative models by expanding the search space to include multiple hypotheses at each generation step. Previous research Yan et al. (2021) has demonstrated that applying beam search in an entity-centric generation does not improve the model’s performance or even degrade it. We conduct experiments on UniNER and our model, the results of which are shown in Table 7. We discover that as the beam size increases, the performance of the UniNER model decreases. In contrast, we observe a consistent improvement with beam search under our task schema. Upon a detailed case study of the model’s generated results, we found that our task schema possesses a self-correction mechanism. The model retains some other hypotheses while generating subsequent results. In the decoding process that follows, the model can correct earlier mistakes. As demonstrated in Fig. 6, the model revises its previous incorrect prediction of “O” for “the fog” upon encountering the subsequent token “rated”. This token is crucial for identifying the entity type associated with “the fog”. A detailed analysis is provided in Appendix E.
@@ 7 Conclusion
This paper explores the potential of a strong Generative Named Entity Recognition (NER) system based on pre-trained LLMs by integrating the negative instances into training. Through experiments, we have demonstrated significant advancements. Our approach, which combines the inclusion of contextual information and a clear definition of entity boundaries through negative instances, has proven to be highly effective in improving the model’s performance, especially in zero-shot scenarios where prediction uncertainty is high. The introduction of an LCS Matching algorithm further addresses the challenges of converting unstructured text into structured entities, ensuring accurate categorization and alignment. These findings highlight the crucial role of negative instances in NER tasks and the potential of generative models to revolutionize the field.
@@ 8 Limitation
Acknowledgements
We want to thank all the anonymous reviewers for their valuable comments. This work was supported by the National Science Foundation of China (NSFC No. 62206194), the Natural Science Foundation of Jiangsu Province, China (Grant No. BK20220488), Young Elite Scientists Sponsorship Program by CAST (2023QNRC001), and Supercomputing Center in Yancheng, Grant No. 20231001.
@@ Appendix B Hyper-parameters settings
GNER LLaMA Omission (39%) 1 raw who directed the film the lorax
pred. who directed the lorax
2 raw any reasonably priced indian restaurants in the theater district
pred. any reasonably priced indian restaurants in theater district
Addition (3%) 3 raw the conservative regionalist navarra suma finished first and  …
pred. the conservative regionalist regionalist navarra suma finished first and  …
Substitution (58%) 4 raw which five star italian restaurants in manattan have the best reviews
pred. which five star italian restaurants in manhattan have the best reviews
5 raw polyethylene terephthalate ( pet ) bottles are made from ethylene and p-xylene .
pred. polyethylene terephthalate ( p e t) bottles are made from ethylene and p-xylene .
GNER T5 Omission (23%) 6 raw …  whose debut album tol cormpt norz norz norz rock hard journalist wolf-rüdiger mühlmann considers a part of war metal ’s roots .
pred. …  whose debut album tol cormpt norz norz rock hard journalist wolf-rüdiger mühlmann considers a part of war metal ’s roots .
7 raw jennifer lien starred in this action film of the the last six years that received a really good rating
pred. jennifer lien starred in this action film of the last six years that received a really good rating
Addition (2%) 8 raw …  performed by wet wet wet that remained at number 1  …
pred. …  performed by wet wet wet wet that remained at number 1  …
9 raw …  liked by many people that starred william forsythe
pred. …  liked by many people that starred william forsythe the
Substitution (75%) 10 raw four more children followed : charlotte brontë , ( 1816-1855 ) , branwell brontë ( 1817-1848 ) , emily brontë , ( 1818-1848 ) and anne ( 1820-1849 ) .
pred. four more children followed : charlotte bront , ( 1816-1855 ) , branwell bront ( 1817-1848 ) , emily bront , ( 1818-1848 ) and anne ( 1820-1849 ) .
@@ Appendix C Problems in long sequence
Appendix D Optimization in LCS Matching
Optimization in Matching Algorithm
We optimize the complexity of the LCS algorithm using a hierarchical divide-and-conquer approach through the following steps: (1) If the sequence does not have the above problem, i.e., $\tilde{X}=X$, it is obvious that $\hat{Y}=\tilde{Y}$. The time complexity is $O(N)$, (2) For the omission case, where $\tilde{X}$ is a subsequence of $X$, the matching process can be accomplished in $O(N)$ through greedy matching. (3) In other cases, we have implemented a fast version of the LCS algorithm Hunt and Szymanski (1977) within $O(N\log N)$, based on the nature of the small number of duplicate words in $\tilde{X}$. Our experimental results in Table 6 demonstrate that the optimization can significantly enhance efficiency, achieving up to a 17.3 times speedup for long sequences compared to the naive $O(N^{2})$ implementation.
Back Tokenization
One notable problem in the matching process is the missing words in the vocabulary, as detailed in our case study in Appendix C. For example, “antropología” in the original text becomes ‘antropologa’ in the model’s predictions, resulting in an inaccurate match in the matching process. To address this, we employ back tokenization, which involves tokenizing each word in the original text and then detokenizing it to match a word in the model’s vocabulary, thereby creating a more resilient matching condition.
Appendix E Self-correction Mechanism
@@
Appendix F Detailed Evaluation Results
@@
GNER LLaMA 1 w/o beam search who(O) is(O) directing(O) the(O) hobbit(B-title)
w/ beam search who(O) is(O) directing(O) the(B-title) hobbit(I-title)
2 w/o beam search what(O) is(O) the(O) plot(O) of(O) the(O) wild(B-title) bunch(I-title)
w/ beam search what(O) is(O) the(O) plot(O) of(O) the(B-title) wild(I-title) bunch(I-title)
3 w/o beam search was(O) there(O) a(O) romantic(O) film(O) noir(O)
w/ beam search was(O) there(O) a(O) romantic(B-genre) film(I-genre) noir(I-genre)
4 w/o beam search does(O) paymon(B-Restaurant Name) serves(O) white(B-Cuisine) wine(I-Cuisine)
w/ beam search does(O) paymon(B-Restaurant Name) serves(O) white(B-Dish) wine(I-Dish)
GNER T5 5 w/o beam search …  some(O) batman(B-character) movies(O) from(O) the(O) 1990s(B-year)
w/ beam search …  some(O) batman(B-title) movies(I-title) from(O) the(O) 1990s(B-year)
6 w/o beam search where(O) was(O) the(O) presidio(B-title) filmed(O)
w/ beam search where(O) was(O) the(B-title) presidio(I-title) filmed(O)
7 w/o beam search …  the(O) third(O) harry(O) potter(O) movie(O) called(O)
w/ beam search …  the(O) third(B-title) harry(I-title) potter(I-title) movie(I-title) called(O)
@@
Appendix G Environmental Impact